# SADG: Segment Any Dynamic Gaussian Without Object Trackers

### v1

@@ Abstract
Understanding dynamic 3D scenes is fundamental for various applications, including extended reality (XR) and autonomous driving. Effectively integrating semantic information into 3D reconstruction enables holistic representation that opens opportunities for immersive and interactive applications. We introduce SADG, Segment Any Dynamic Gaussian Without Object Trackers, a novel approach that combines dynamic Gaussian Splatting representation and semantic information without reliance on object IDs. In contrast to existing works, we do not rely on supervision based on object identities to enable consistent segmentation of dynamic 3D objects. To this end, we propose to learn semantically-aware features by leveraging masks generated from the Segment Anything Model (SAM) and utilizing our novel contrastive learning objective based on hard pixel mining. The learned Gaussian features can be effectively clustered without further post-processing. This enables fast computation for further object-level editing, such as object removal, composition, and style transfer by manipulating the Gaussians in the scene. We further extend several dynamic novel-view datasets with segmentation benchmarks to enable testing of learned feature fields from unseen viewpoints. We evaluate SADG on proposed benchmarks and demonstrate the superior performance of our approach in segmenting objects within dynamic scenes along with its effectiveness for further downstream editing tasks. Our project page is available at: https://yunjinli.github.io/project-sadg/.
@@ 1 Introduction
We live in a dynamic, three-dimensional world. Most of the time, we describe our environment at a high level of objects. As we move around and observe the scene from different viewpoints, we complete the geometry and perceive the colors contributing to our object representation. At the same time, we infer dynamic properties with object motion and deformations. Despite pose, color, and shape changes, the underlying semantic meaning stays constant. This intuition can serve as a strong constraint for neural 3D representations that have gained recent popularity and promote innovation in augmented reality, gaming, and autonomous driving interactive applications. In these technologies, seamlessly integrating semantic information into underlying representation ensures an immersive experience and gives users advanced control over the dynamic scene contents.
Neural Radiance Fields (NeRFs) have revolutionized the field of 3D reconstruction and demonstrated excellent performance in synthesizing photorealistic novel views of a scene from any angle. The limitations of vanilla NeRFs promoted many follow-up works, including those that encompass semantic information by lifting 2D segmentation masks [24, 7, 19] or distilling features [21, 18] and those that faithfully represent dynamic scenes [41, 35, 27]. Due to continuous implicit representation, these methods are computationally intensive as they require re-rendering from multiple views to ensure edit consistency, which hinders their applicability for interactive downstream tasks.
Building on the strengths of NeRFs, Gaussian splatting (3DGS) [17] has introduced a powerful new approach that prioritizes efficiency and speed. Dynamic versions of 3DGS [30, 56] have offered faithful reconstruction of scenes with motion. These models can serve as a geometric prior for learning semantic features that are consistent in both space and time and open new possibilities for real-time interactive applications like scene editing and dynamic interactions. Towards this, a line of works has extended geometric properties of colored and explicit Gaussians with semantic information [6, 57], producing consistent novel view synthesis and segmentation. Nonetheless, these works are limited to static scenes and cannot handle objects in motion. There are only a few concurrent works [16, 25], which unify semantics and dynamics to comprehensively understand real-world environments. However, SA4D [16] relies on the supervision of object trackers and suffers from semantic identity conflicts in multi-view scenes. Moreover, their method is sensitive to mask noise and requires post-processing to remove degenerate Gaussians. At the same time, DGD [25] has a higher memory requirement and longer training time as they are learning 4D feature fields based on 512-dimensional CLIP [42] or 384-dimensional DINOv2 [34]. Our work aims to tackle both single- and multi-view representations with fast rendering times, minimal post-processing, and low memory footprint.
To this end, we introduce our novel framework, Segment Any Dynamic Gaussian Without Object Trackers (SADG), which effectively combines dynamic Gaussian splatting backbone [56] and semantic information without any reliance on trackers that maintain consistent object identities. To learn the spatio-temporal semantic field, we define a new contrastive objective based on hard positive and negative mining with the Segment Anything Masks (SAM) [20]. We inherit the strengths of zero-shot 2D scene understanding of the foundation model to learn the semantically-aware feature field and produce consistent spatio-temporal segmentation. Our 32-dimensional features are compact and do not depend on features of a 2D segmentation model. Rendering time and storage requirements are almost unaffected, enabling real-time interaction and editing. The learned feature field can be effectively clustered without any further post-processing and utilized in a range of downstream editing tasks, including object style transfer [14], re-coloring, composition, and removal. Given an early stage of development and benchmarks for semantic segmentation for dynamic scenes, we propose to extend existing novel-view synthesis datasets with semantic benchmarks. The proposed evaluation protocol includes single- and multi-view scenes with challenging motions and diverse contents, allowing extensive testing of semantically-aware latent representations from unseen viewpoints.
@@
We propose Segment Any Dynamic Gaussian Without Object Trackers (SADG), a new approach that achieves multi-view consistent segmentation of dynamic scenes without any tracking supervision.
We leverage information from 2D masks to learn semantically-aware latent representation for dynamic scenes using a novel contrastive learning objective.
We extensively evaluate our method on single- and multi-view scenes across five dynamic novel-view benchmarks and demonstrate state-of-the-art segmentation performance.
We demonstrate the general nature of our feature space and apply it to several downstream tasks, including object removal, style transfer, and scene composition.
We offer an interactive framework that allows to edit the scene with a simple mouse click or a text prompt. Our tool does not depend on input IDs or 2D strokes and operates in real-time.
@@ 2 Related Works
2.1 Semantic Feature Learning
With the emergence of various vision foundation models, such as Segment Anything Model (SAM) [20], DINO [5], and DINOv2 [34], researchers start to integrate the semantics from these foundation models into their reconstruction. SA3D [7] is the pioneering method that extends 2D SAM masks across images into 3D consistent object masks. GARField [19] introduces hierarchical grouping, leveraging the physical scales of 2D masks to group objects of different sizes and represent scenes at different granularity scales. Gaussian Grouping [57] is a pioneering Gaussian-based approach that utilizes consistent object IDs generated by video tracker DEVA [8] to render object IDs into camera views. Concurrent work such as SAGA [6] utilizes the masks generated from SAM, combined with scales obtained from 3D data inspired by GARField, to segment objects of various sizes. Notably, these methods are designed for static scenes. To date, only two concurrent work, SA4D [16] and DGD [25], addresses the segmentation task in dynamic 3DGS. DGD extends the rasterization pipeline of 3DGS to be able to render semantic feature of each Gaussian. The rendered Gaussian feature map is then trained with the supervision of DINOv2 [34] or CLIP [42] to mimic the output from their feature extractors. The biggest drawback of such supervision is anticipated longer training times as the dimensions of the features generated by DINOv2 and CLIP are quite large (384 and 512). Similar to Gaussian Grouping and its derivatives [11, 31], SA4D relies on video trackers to associate masks across different views to produce consistent object IDs. These approaches face significant limitations, as the mask association often fails to offer consistent object IDs in dynamic scenes, such as the oven-mitts sequence from HyperNeRF [36] as shown in Fig. 2.
@@ 2.2 Scene Editing Applications
Learning semantically-aware latent space opens opportunities for interactive downstream applications, allowing users to manipulate scenes on an object level. Several works proposed editing techniques for radiance field representations [13, 49], which are limited to a single object. On the other hand, scene-based methods [53, 24] learn a separate radiance field per object and render the full scene by combining learned opacities along the viewing rays. While the structure is modular, its editing application is limited to a pre-defined number of objects. Our method is guided by user interaction and is capable of operating on any object in the scene. Another line of work distills features such as DINO [34] into volumetric representation and shows editing capabilities on learned continuous feature field [48]. Due to the implicit representation, scene editing requires a re-assessment of the final consistency by rendering from novel scenes, which impedes the model’s applicability to real-time interactive applications.
At the same time, Gaussian splatting methods mitigate the runtime limitation of NeRFs with their explicit point-based representation. SAGA [57] proposes Local Gaussian Editing scheme, which localize the Gaussians and manipulate them according to the editing task. Feature3DGS [60] distills features from 2D foundation models such as SAM [43] and LSeg [26] enabling promptable and language-driven 3D edits. Nonetheless, these methods operate only on static scenes, while our method is capable of editing dynamic objects and ensuring temporal consistency of the edits. DGD [25] shows semantic texture editing performance by utilizing a pre-trained diffusion model and fine-tuning dynamic 3DGS representation with the reconstruction loss. On the other hand, we demonstrate a variety of editing tasks, including style transfer, object removal, and composition. Despite offering several scene editing opportunities, SA4D [16] has limited user interactivity and requires prior knowledge of object identities. Our framework contains a user-friendly graphical interface that allows operation with a simple text prompt or a mouse click.
@@ 3 Method
In this section, we introduce our novel framework SADG. As illustrated in Fig. 1 the pipeline comprises two main components: dynamic geometry reconstruction (Sec. 3.1) and Gaussian feature learning (Sec. 3.2). We adopt the Deformable-3DGS [56] pipeline to reconstruct the 4D scene. Once the 4D reconstruction is learned, we freeze this representation and proceed with Gaussian feature learning using SAM [20] masks and our novel contrastive learning objective in the rendered feature space.
@@ 3.1 Dynamic Geometry Reconstruction
We follow the approach proposed in [56] and describe it briefly for completeness. Unlike Dynamic3DGS [30] model, where the reconstruction is stored per frame, resulting in high memory consumption, an MLP learns per-Gaussian deformation with respect to the static canonical space , which is defined as
is the center position of the i-th Gaussian, is the quaternion representing the rotation, is the scaling vector, refers to the opacity, and is the Spherical Harmonic Coefficients encoding the color information () as per [56].
@@ 3.2 Gaussian Feature Learning
Problem statement Most of the existing works [16, 57] utilize a video tracker to provide consistent object mask IDs across views and render the segmentation field directly by predicting the per-pixel probability of different object classes. However, such approaches have a significant bottleneck, namely their reliance on perfect object mask association across views. When mask inconsistencies of the same object occur between views, the entire optimization pipeline can break down, leading to sub-optimal results or failure.
Therefore, we propose a method that aims to learn the 3D semantics of objects and leverage them directly for downstream tasks across different views. In other words, instead of predicting per-pixel probabilities during rendering, we focus on integrating semantic information into Gaussians and leverage semantic constancy for learning multi-view consistent spatio-temporal feature field.
Representation Given 4D reconstruction, we incorporate semantic information into the scene. Specifically, we extend its -th Gaussian with a 32-dimensional semantic feature vector , the so-called “Gaussian feature”. The semantically-aware feature field of a scene represented with Gaussians can be expressed as
Rendering Gaussian Features The learned Gaussian features can be rendered similarly to [17, 60] with point-based -rendering [23, 22]. Given a sorted subset of Gaussians that are related to pixel , the resulting rendered features can be computed as follows:
Existing methods [6] use smooth Gaussian features instead, as they help to filter out noisy 3D Gaussian features with high similarity scores corresponding to unrelated objects. The smooth Gaussian feature for the -th Gaussian is computed as per Eq. 6, where denotes a set of nearest neighbors [9] of the -th Gaussian based on its 3D position. We adopt this convention in our work and utilize smooth feature vectors in the rendering process (Eq. 5).
Our contrastive semantically-aware learning objective
The generated set of SAM masks comprise binary masks of a given image , where each mask corresponds to a different object. In practice, we do not consider all masks from SAM to alleviate memory and computational requirements. Instead, we randomly sample masks for further usage.
@@
We define the pixel-mask correspondence vector in Eq. 7, which captures mask information for a certain pixel coordinate (, ) across all binary masks in . This vector is similar to a one-hot encoding technique; however, there is no guarantee that a pixel belongs to a single mask due to 2D segmentation inaccuracies. We visually demonstrate the formulation of in Fig. 4.
where and .
In practice, we only sample pixels from an image such that . Mathematically, we define the sampled pixel coordinates as a set .
@@
The pairwise similarities of the sampled pixel-mask correspondence vectors can be retrieved by computing the Gram matrix of their stacked matrix . All entries of the resulting Gram matrix are then bounded to either 0 (negative pair) or 1 (positive pair) to form a mask-based correspondence matrix among sampled pixels , formally defined as
Additionally, we formulate a correspondence matrix for the sampled rendered features . Analogous to the mask-based correspondence matrix , we compute a feature-based Gram matrix, capturing the pairwise similarities between the features as follows:
where is defined as follows:
Finally, we have all the necessary ingredients to formulate the contrastive learning objective based on a hard positive and negative mining scheme. Guided by the mask-based correspondence matrix, we enforce features to be close for pixels from the same object mask and far apart for the distinct objects. Furthermore, we concentrate on hard pixel pairs. Specifically, we sample positive samples with low similarity scores and negative samples with high similarity scores. With this, we formally define the optimization objective as
Note that is an indicator function. It returns 1 when , and 0 otherwise. stands for hard positive which is defined as and refers to a hard negative which is defined as . Here, and are the positive and negative thresholds for pairwise similarity respectively. As both and are symmetric matrices following the property of the Gram matrix, we only consider entries in the upper triangular part and exclude the diagonal to avoid repeated computation. We show rendered Gaussian features learned with our contrastive scheme in Fig. 1. For demonstration, they are compressed into via PCA [33], allowing each rendered feature to be represented with RGB color.
@@
Gaussian Feature Clustering To enable interactive applications with the learned 4D representation, we deploy the DBSCAN algorithm [12] to cluster the Gaussian features into several groups and generate the so-called Gaussian Clusters as defined in Eq. 12. provides the corresponding cluster ID for -th Gaussian feature:
Note that refers to the total number of clusters. We demonstrate them visually in Fig. 1. As can be observed, they represent distinct objects in 4D space.
Since a typical scene contains a large number of Gaussians (more than 300k), we adopt the sub-sampling scheme proposed by [6], where only 2% of the Gaussian features are used for clustering. We propagate the cluster IDs to the remaining of the Gaussians by computing the feature similarity between cluster-representative Gaussians. In the supplementary material, we demonstrate that the performance is not affected by a significant reduction in the number of Gaussians contributing to the generation of clusters.

### v2

# TRASE: Tracking-free 4D Segmentation and Editing

@@ Abstract
Understanding dynamic 3D scenes is crucial for extended reality (XR) and autonomous driving. Incorporating semantic information into 3D reconstruction enables holistic scene representations, unlocking immersive and interactive applications. To this end, we introduce TRASE, a novel tracking-free 4D segmentation method for dynamic scene understanding. TRASE learns a 4D segmentation feature field in a weakly-supervised manner, leveraging a soft-mined contrastive learning objective guided by SAM masks. The resulting feature space is semantically coherent and well-separated, and final object-level segmentation is obtained via unsupervised clustering. This enables fast editing, such as object removal, composition, and style transfer, by directly manipulating the scene’s Gaussians. We evaluate TRASE on five dynamic benchmarks, demonstrating state-of-the-art segmentation performance from unseen viewpoints and its effectiveness across various interactive editing tasks. Our project page is available at: https://yunjinli.github.io/project-sadg/
@@ 1 Introduction
Humans experience the world in motion, where objects persist over time despite changes in pose, shape, and appearance. Our ability to recognize objects remains stable, even as we move and observe them from different vantage points. This phenomenon suggests the existence of neural 3D representations, which preserve semantic consistency across both space and time. Such representations are central for applications in augmented reality, gaming, and autonomous systems, where interactive manipulation of dynamic scenes requires object-level segmentation that is spatio-temporally consistent and efficient.
Neural Radiance Fields (NeRFs) [31] have significantly advanced 3D reconstruction by enabling photorealistic novel view synthesis, but their implicit nature makes them computationally too expensive for real-time interaction. Gaussian Splatting (3DGS) [20] has recently emerged as a powerful alternative, achieving real-time rendering with explicit and compact scene representations. Dynamic extensions of 3DGS [28, 52] allow faithful reconstruction of moving scenes, but they lack built-in object segmentation and rely purely on geometry and color. While several works have integrated semantic information into static 3D Gaussian representations [6, 53], they fail to handle motion, deformation, and viewpoint variation in dynamic scenes.
Only a handful of works have explored the unification of semantics and dynamics in 3DGS [19, 24]. SA4D [19] incorporates segmentation into dynamic Gaussian splatting, using video object trackers to match masks across views similar to Gaussian Grouping [53]. However, approaches using these trackers often suffer from identity switches, causing inconsistent segmentation in multi-view settings - particularly in the presence of occlusions, non-rigid deformations, or rapid motion, as shown in Fig. 3. DGD [24] addresses temporal consistency by distilling semantically-aware features from foundation models like CLIP [45] and DINOv2 [32], enabling tracking. Yet, these features are high-dimensional, leading to significant computational overhead from both the large memory requirements and slow training.
In this work, we present TRASE, a tracking-free segmentation framework for dynamic scene understanding. TRASE learns temporally consistent segmentation features in a contrastive setting using only 2D binary masks. Unlike prior contrastive segmentation methods [12, 13, 54], which explicitly enforce cluster formation during training and promote statistical consistency of features, TRASE applies contrastive learning directly at the pixel-pair level in a fine-grained rendered feature space, deferring object clustering to a final post-training step. This design enables TRASE to learn a segmentation field with both high temporal consistency and sharp, well-defined object boundaries. It also effectively mitigates common artifacts, such as floaters and out-of-FoV errors, that frequently degrade prior contrastive learning approaches.
To the best of our knowledge, we are the first to introduce a benchmark for segmentation in novel views of dynamic scenes. While established datasets like DAVIS [35, 36, 3, 4] and VOS [47, 50, 51] address dynamic video segmentation, none evaluate segmentation quality in unseen viewpoints. Moreover, our benchmark is significantly larger than the evaluation protocol suggested in [24], as our test sequences comprise single- and multi-view scenes and captures real-world scenarios. Beyond achieving state-of-the-art segmentation accuracy, TRASE is tailored for real-time interactive scene editing. Our compact 32-dimensional feature space requires minimal post-processing and integrates seamlessly into object manipulation tasks such as style transfer, recoloring, composition, and removal.
@@
We propose TRASE, a new approach for multi-view consistent segmentation of dynamic scenes without tracking labels. To achieve this, we design a novel contrastive learning objective based on differentiable feature rendering and contrastive learning, ensuring segmentation stability across time and views.
We offer a broad benchmark for segmentation in novel views of dynamic scenes, which encompass a broad spectrum of camera setups and captured scenarios.
We extensively evaluate our method on five dynamic datasets and achieve state-of-the-art segmentation performance in both single- and multi-view settings.
We demonstrate the applicability of our feature space to several scene editing tasks, including object removal, style transfer, and scene composition.
@@ 2 Related Works
2.1 3D Segmentation
In static scenes. With the emergence of various vision foundation models, such as Segment Anything Model (SAM) [22], DINO [5], and DINOv2 [32], researchers start to integrate the semantics from these foundation models into their reconstruction. SA3D [7] is the pioneering method that extends 2D SAM masks across images into 3D consistent object masks. GARField [21] introduces hierarchical grouping, leveraging the physical scales of 2D masks to group objects of different sizes and represent scenes at different granularity scales. Gaussian Grouping [53] first proposes to utilize consistent object IDs generated by video tracker DEVA [8] to render object IDs into camera views. Concurrent work such as SAGA [6] utilizes the masks generated from SAM, combined with scales obtained from 3D data inspired by GARField, to segment objects of various sizes. GAGA [29] proposes a mask group identity assignment technique using a 3D-aware memory bank to compensate for inaccurate semantic labels and use group IDs as pseudo-labels for identity encoding. Our model implicitly learns multi-view consistent semantic Gaussian features using our designed contrastive loss in a single training stage without identity assignments.
In dynamic scenes. To the best of our knowledge, only two prior works, SA4D [19], and DGD [24], address the segmentation task in dynamic 3DGS. DGD extends the rasterization pipeline of 3DGS to be able to render the semantic features of each Gaussian. The rendered Gaussian feature map is then trained with the supervision of DINOv2 [32] or CLIP [37] to mimic the output from their feature extractors. The drawbacks of such supervision are longer training times and bigger 3DGS models, as the dimensions of the features generated by DINOv2 and CLIP are quite large (384 and 512), and the Gaussian features are also stored. Following the identity encoding concept from Gaussian Grouping and its derivatives [10, 29], SA4D introduces a temporal identity feature field. As the method relies on consistent object IDs for supervision, it is sensitive to temporal tracking inaccuracies due to occlusion or fast motion. A concurrent work, Split4D [17], follows several proposed design choices, including single-frame supervision and a contrastive learning objective, and utilizes a recent 4D reconstruction model to achieve 4D segmentation. Our TRASE does not rely on tracking labels and effectively handles single- and multi-view data with fast rendering times, minimal post-processing, and low memory footprint.
2.2 Contrastive Learning for 3D Segmentation
A number of recent works define multi-view 3D segmentation as a contrastive learning objective. Contrastive Gaussian Clustering (CGC) [12] training aims to maximize intra-cluster similarity in the 2D rendered feature space by enforcing the feature’s proximity with its corresponding cluster’s mean feature and ensuring its spatial consistency. OmniSeg3D [54] follows a similar CGC objective and supervises 3D point clustering using SAM-based image patches. ContrastiveLift [13] proposes a slow-fast contrastive fusion strategy, where a slow network aggregates global object-level information while a fast network refines per-point features. OpenGaussian [46] enforces intra-cluster similarity and inter-cluster dissimilarity of rendered features and further trains their discretized version. Split4D [17] utilizes the same objective as [12] and additionally enforces the features to align with DINOv2 [32] in the early training stage. Our TRASE proposes a contrastive objective based on soft sample mining of pixel pairs, which focuses on only imperfect features and allows us to learn the segmentation field effectively in 4D. No costly distillation is required as we supervise the training with binary 2D masks.
@@ 2.3 Scene Editing Applications
Learning a semantically aware latent space enables interactive scene editing at the object level. Existing radiance-field editing methods are either restricted to single objects [16, 45] or rely on per-object scene decompositions [49, 23], which limit editing to a fixed set of instances. Other approaches, such as DINO [32], distill features into volumetric fields [44], but their implicit representations require expensive re-rendering for consistency, hindering real-time interaction. Gaussian-based methods [20] alleviate NeRF runtime issues via explicit point-based primitives: SAGA [53] enables local Gaussian editing, and Feature3DGS [56] leverages 2D foundation models like SAM [38] and LSeg [25] for promptable, language-guided 3D edits, yet all focus on static scenes. In contrast, our method edits dynamic objects with temporally consistent results and supports diverse tasks, including style transfer, object removal, and composition [24]. While SA4D [19] offers semantic scene editing, it requires prior object labels and limited interaction. In contrast, our framework provides a user-friendly graphical interface that supports simple text prompts and mouse-based selection.
@@ 3 Method
In this section, we introduce our novel TRASE. As illustrated in Fig. 2, the pipeline comprises two main components: dynamic geometry reconstruction (Sec. 3.1) and Gaussian feature learning (Sec. 3.2). We adopt the Deformable-3DGS [52] pipeline to reconstruct the 4D scene. Once the 4D reconstruction is learned, we freeze this representation and proceed with Gaussian feature learning using SAM [22] masks and our novel contrastive learning objective in the rendered feature space.
@@ 3.1 Dynamic Geometry Reconstruction
We follow the approach proposed in [52] and describe it briefly for completeness. Unlike Dynamic3DGS [28] model, where the reconstruction is stored per frame, resulting in high memory consumption, an MLP learns per-Gaussian deformation with respect to the static canonical space , which is defined as for . is the center position of the i-th Gaussian, is the quaternion representing the rotation, is the scaling vector, refers to the opacity, and is the Spherical Harmonic Coefficients encoding the color information () as per [52].
@@ 3.2 Gaussian Feature Learning
Summary. Our feature learning augments each 3D Gaussian with a 32-dimensional feature vector, enabling segmentation without explicit tracking labels. It is worth noting that, in contrast to view-dependent spherical encoding, our semantic feature vector is consistent across different views and time. Thanks to differentiable feature rendering [56] and SAM [22] masks, we construct and explicitly enforce pixel-mask associations in the Gaussian feature field. Our novel contrastive learning objective is designed to be efficient by sub-sampling and effective by integrating SAM [22] masks. Using only 2D guidance and without any object labels, Gaussians are trained to separate per-pixel rendered features from different segments and align those within the same segment. To achieve this, we design a soft-mining objective. Unlike hard mining, which selects pairs strictly by similarity, soft mining uses a relaxed, existential rule that increases pixel coverage and makes training less sensitive to noisy feature similarities. Finally, the learned features are globally clustered, producing multi-view consistent segmentation.
@@
Our notation. A generated set of SAM masks comprise binary masks of a given image , where each mask corresponds to a different segment. With this, we define the pixel-mask correspondence vector , which captures information across all binary masks in for a certain pixel coordinate (, ). This vector is similar to a one-hot encoding technique; however, there is no guarantee that a pixel belongs to a single mask due to 2D segmentation inaccuracies. We visually demonstrate the formulation of in Fig. 4.
@@
The similarities of pixel-mask vectors for a pair of pixels (, ) can be retrieved by computing the Gram matrix of their stacked matrix , where refers to the number of sampled pixels. All entries of the resulting Gram matrix are then bounded to either 0 (negative pair) or 1 (positive pair) to form a mask-based correspondence matrix among sampled pixels, formally defined as
Analogous to the binary mask-based Gram matrix , we compute a feature-based matrix , capturing the pairwise similarities between the rendered features . Following common practices [6], we consider a smooth version of Gaussian features, as it is more robust than considering individual features, which might have high similarity scores corresponding to unrelated objects. The smooth Gaussian feature is computed as the mean feature of its nearest neighbors [9] based on the Gaussian’s 3D position.
Our learning objective. To robustly learn multi-view coherent feature embeddings, we adopt a soft-mined contrastive learning objective. Unlike standard hard mining, which selects positive or negative pairs strictly based on their individual similarity values, soft mining employs an existential criterion: a positive pixel pair (,) is considered into final positive loss computation if there exists any positive pair involving whose similarity falls below the positive threshold . Similarly, a pair is included in the negative loss if there exists any negative pair involving whose similarity exceeds the negative threshold . By design, the soft-mined strategy broadens pixel coverage during training, in contrast to naive hard sampling that typically concentrates on a sparse set of imperfect pixel pairs and leaves many pairs untrained. We further restrict the selection to non-redundant pairs by considering only the upper-triangular portion of the matrices (). Formally, the soft-mined masks are defined as
The positive and negative losses are then computed as weighted sums over these masks:
where is a normalization factor (e.g., the total number of pixel pairs), is a pixel-weight matrix similar to [6], and ensures that only positive contributions are considered for negative pairs. The final loss is then defined in Eq. 8 with an extra regularization term to stabilize the feature learning.
@@
Inference. To enable interactive applications with the learned 4D representation, we deploy DBSCAN algorithm [11] to cluster the Gaussian features into several groups and generate the so-called Gaussian clusters. In contrast to the K-Means algorithm [27, 30], it does not require any prior knowledge of the number of clusters in a scene. As can be seen in Fig. 2, clusters represent distinct objects, which verifies accurate alignment of our learned feature space with the object-level semantic field.
We further observe that the object boundaries can be effectively refined by applying a simple similarity threshold, which prunes Gaussians with features that are less similar to the corresponding cluster’s mean. This simple technique enhances the method’s performance in challenging segmentation scenarios, such as those involving occlusions and affected by sparse observations.